OpenStudy (anonymous):

√49/√2 1) √2/2 2) 7/√2 3) 7√2/2 4) 7/2

OpenStudy (anonymous):

2) 7/√2

OpenStudy (anonymous):

because √49 = 7

OpenStudy (anonymous):

49 = 7^2

OpenStudy (anonymous):

\[\sqrt{49}=7\] So \[\frac{7}{\sqrt{2}}\]